# A Survey on Statistical Theory of Deep Learning: Approximation, Training Dynamics, and Generative Models

### v1

@@ 2.2 Non-parametric Regression via fully-connected networks
Several remarks follow in the sequel. First, the rate $\phi_{n}$ is dependent on the effective dimension $t_{i}$ which can be much less than the ambient dimension $d_{0}$, and this implies that the neural net can circumvent the curse of dimensionality. Second, the depth $L$ should be chosen to scale with the sample size in a rate $\log n$. Third, the width $\mathbf{p}_{\text{max}}$ can be chosen to be independent of the smoothness indices. Lastly, the result is saying what is important for the statistical performance is not the sample size, but the amount of regularization. This is explicitly reflected through the number of active parameters $\mathcal{N}\asymp n\phi_{n}\log n\ll n$.
Schmidt-Hieber (Thm 5) Lu et al. (Thm 1.1.) Jiao et al. (Corol. 3.1)
Depth $\mathcal{O}(\lceil\log_{2}(r\vee d)\rceil)$ $\mathcal{O}(r^{2}+d)$ $\mathcal{O}((\lfloor r\rfloor+1)^{2})$
Width $\mathcal{O}((d+\lceil r\rceil)(r+1)^{d})$ $\mathcal{O}(r^{d+1}d3^{d})$ $\mathcal{O}((\lfloor r\rfloor+1)^{2}3^{d}d^{\lfloor r\rfloor+1})$
Approx. Rate $\mathcal{O}(6^{d}(r+1)^{d})$ $\mathcal{O}(8^{r}(r+1)^{d})$ $\mathcal{O}((\lfloor r\rfloor+1)^{2}d^{\lfloor r\rfloor+(r\vee 1)/2})$
Table 1: A summary of $r$ and $d$ dependences in the prefactors of constructed deep ReLU networks’ depths, widths and approximation rates for approximating the functions $f\in W^{r}_{\infty}(\mathcal{C}^{d})$ in Schmidt-Hieber (2020); Lu et al. (2021); Jiao et al. (2021).
Explicit dependence on $d$ in the pre-factor Csuperscriptnormal-$C^{\prime}$. After the publication of Schmidt-Hieber (2020), a series of works further studied the approximation and estimation of $f\in W^{r}_{\infty}(\mathcal{C}^{d})$ via deep ReLU FNNs, where $\mathcal{C}^{d}$ is a $d$-dimensional cube. Specifically, Ghorbani et al. (2020a) pointed out that the additive function studied in Schmidt-Hieber (2020) has an exponential dependence in $d$ in the prefactor $C^{\prime}$ of the estimation error, requiring $n\succsim d^{d}$ sample sizes for good convergence rate. Later,  Shen et al. (2021) and Lu et al. (2021) tracked the explicit dependence on $d$ in the approximation error as well as in the architectural components, (i.e., width ($\mathcal{W}$) and depth ($L$)), for approximating $f\in C(\mathcal{C}^{d})$ (i.e., Lipschitz continuous functions on $\mathcal{C}^{d}$) and $f\in W^{r}_{\infty}(\mathcal{C}^{d})$, respectively. Specifically, Lu et al. (2021) improved the prefactor $\mathcal{O}(8^{r}(r+1)^{d})$ when it is compared with $\mathcal{O}(6^{d}(r+1)^{d})$ from Schmidt-Hieber (2020) for the approximation error. Most recently, Jiao et al. (2021) further improved the prefactor $\mathcal{O}((\lfloor r\rfloor+1)^{2}d^{\lfloor r\rfloor+(r\vee 1)/2})$ for approximating $f\in W^{r}_{\infty}(\mathcal{C}^{d})$. In Table 1, comparisons of depths, widths, and approximation rates of the constructed deep ReLU networks in Schmidt-Hieber (2020); Lu et al. (2021); Jiao et al. (2021) are summarized. Along this line of research, Suh et al. (2022) showed the deep ReLU net can avoid the curse when the smoothness index $r$ grows in the order of data dimension $d$, i.e., $r=\mathcal{O}(d)$, requiring only $\mathcal{O}(d^{2})$ number of active parameters for approximating functions in Hölder spaces defined on a unit sphere.
Approximation of $f\in W_{\infty}^{r}(\mathcal{C}^{d})$ via deep ReLU FNN. Approximation theory of deep ReLU FNN for functions $f\in W_{\infty}^{r}(\mathcal{C}^{d})$ has a lengthy history in the literature. Representatively, Mhaskar (1996) showed that $f$ can be approximated uniformly within $\varepsilon$-approximation accuracy with a $1$-layer neural network of $\mathcal{O}(\varepsilon^{-d/r})$ neurons and an infinitely differentiable activation function. Later, for deep ReLU networks, Yarotsky (2017) showed that $\mathcal{N}$ in networks is bounded by $\mathcal{O}(\varepsilon^{-d/r}\log\big{(}\frac{1}{\varepsilon}\big{)})$, and the depth has the order $\mathcal{O}(\log(\frac{1}{\varepsilon}))$. He further proved that $\mathcal{N}$ is lower-bounded by the order $\mathcal{O}(\varepsilon^{-d/r})$, which is backed up by the result in DeVore et al. (1989). For $f\in W_{p}^{r}(\mathcal{C}^{d})$ with $1\leq p\leq\infty$Petersen and Voigtlaender (2018) showed that there exists a deep ReLU network with bounded and quantized weight parameters, with $\mathcal{O}(\varepsilon^{-d/r})$ network size, and with $\varepsilon$-independent depth for achieving the $\varepsilon$-accuracy in the $L_{p}$ norm. For approximating functions $f\in W_{\infty}^{r}(\mathcal{C}^{d})$,  Schmidt-Hieber (2020) proved that a network of size $\mathcal{O}(\varepsilon^{-d/r})$ with bounded weight parameters achieves $\varepsilon$-approximation error in the $L_{\infty}$ norm.
Function spaces with special structures. The result of Yarotsky (2017) implies that deep ReLU net cannot escape the curse of dimensionality for approximating $f\in W_{\infty}^{r}(\mathcal{C}^{d})$. Many papers have demonstrated that the effects of dimension can be either avoided or lessened by considering function spaces different from Sobolev spaces, but defined over $\mathcal{C}^{d}$. Just to name a few, Mhaskar et al. (2016) studied that a function with a compositional structure with regularity $r$ can be approximated by a neural network with $\mathcal{O}(\varepsilon^{-2/r})$ neurons within $\varepsilon$ accuracy.  Suzuki (2018) proved the deep ReLU network with $\mathcal{O}(\varepsilon^{-1/r})$ neurons can avoid the curse for approximating functions in mixed smooth Besov spaces.  Chen et al. (2019a) showed the network size scales as $\mathcal{O}(\varepsilon^{-D/r})$ for approximating $C^{r}$ functions, when they are defined on a Riemannian manifold isometrically embedded in $\mathbb{R}^{d}$ with manifold dimension $D$ with $D\ll d$.  Montanelli and Du (2019) and Blanchard and Bennouna (2022) showed respectively the deep and shallow ReLU network break the curse for Korobov spaces.
Estimation rates of excess risk. Many researchers also have tried to tackle how neural networks avoid the curse by considering specially designed function spaces under the non-parametric regression framework. We only provide an incomplete list of them. Such structures include additive ridge functions (Fang and Cheng, 2022), composite function spaces with hierarchical structures (Schmidt-Hieber, 2020; Han et al., 2022), mixed-Besov spaces (Suzuki, 2018), Hölder spaces defined over a lower-dimensional manifold embedded in $\mathbb{R}^{d}$ (Chen et al., 2022a). They all showed the function estimators with neural network architectures can lessen the curse by showing the excess risks of the estimators are bounded by $\mathcal{O}(n^{-2r/(2r+D^{\prime})})$, where $n$ denotes the size of a noisy dataset and $D^{\prime}\ll d$ is an intrinsic dimension uniquely determined through the characteristics of function spaces when they are compared with the minimax risk $\mathcal{O}(n^{-2r/(2r+d)})$ (Donoho and Johnstone, 1998) for $f\in W_{\infty}^{r}(\mathcal{C}^{d})$.
2.3 Non-parametric Classification via fully-connected networks
@@
In this subsection, we consider a binary classification problem via fully-connected networks. Classifiers built with neural networks handle large-scale high-dimensional data, such as facial images from computer vision extremely well, while traditional statistical methods often fail miserably. We will review papers that have attempted to provide theoretical explanations for such empirical successes in high-dimensional classification problems, beyond the existing statistical literature. This is a less studied area than nonparametric regression, although deep learning is often used for classification.
@@
Results on Neural Networks. The table summarizes the settings and results of the works that study the convergence rates of excess risks for binary classification problems of neural networks. Hereafter, $\alpha>0$ indicates the smoothness index of either target function class $(\Sigma)$ in CAR or boundary class $(\mathcal{G})$ in CAD, and $q>0$ denotes the term in the exponent in Margin Assumption (MA).
Reference Function Space Loss Condition Rate
Kim et al. (2021) ReLU FNNs Hinge MA, CAD $\mathcal{O}\bigg{(}n^{-\frac{\alpha(q+1)}{\alpha(q+2)+(d-1)(q+1)}}\bigg{)}$
MA, CAR $\mathcal{O}\bigg{(}n^{-\frac{\alpha(q+1)}{\alpha(q+2)+d}}\bigg{)}$
Feng et al. (2021) ReLU CNNs Hinge CAR $\mathcal{O}\bigg{(}n^{-\frac{r}{2\beta(d-1)+r(2-\tau)}}\bigg{)}$
p-norm $\mathcal{O}\bigg{(}n^{-\frac{pr}{2(\beta+1)(d-1)+2pr(2-\tau)}}\bigg{)}$
2-norm MA, CAR $\mathcal{O}\bigg{(}n^{-\frac{2rq}{(2+q)((\beta+1)(d-1)+2\tau)}}\bigg{)}$
Shen et al. (2022) Hinge MA, CAR $\mathcal{O}\bigg{(}n^{-\frac{r(q+1)}{d+2r(q+1)}}\bigg{)}$
Logistic CAR $\mathcal{O}\bigg{(}n^{-\frac{r}{2d+4r}}\bigg{)}$
Least Square $\mathcal{O}\bigg{(}n^{-\frac{4r}{3d+16r}}\bigg{)}$
Zhou and Huo (2023) ReLU FNNs Hinge
MA:
GMM
$\mathcal{O}\bigg{(}n^{-\frac{q+1}{q+2}}\bigg{)}$
Ko et al. (2023) ReLU FNNs Logistic
MA, CAR:
BA
$\mathcal{O}\bigg{(}n^{-\frac{q+1}{3(q+2)}}\bigg{)}$
Hu et al. (2020) ReLU FNNs 0-1 Loss
MA, CAD:
Teacher-Student
$\mathcal{O}\bigg{(}n^{-\frac{2}{3}}\bigg{)}$
@@
Two recent works studied the binary classification by ReLU convolutional neural networks Feng et al. (2021); Shen et al. (2022). The paper Feng et al. (2021) considered the $p$-norm loss $\phi(t):=\max\{0,1-t\}^{p}$ for $p\geq 1$, and the input data is supported on the sphere $\mathcal{S}^{d-1}$. In the paper, the approximation error bounds and excess risk bounds are derived under a varying power condition, and the target function class is the Sobolev space $W_{p}^{r}(\mathcal{S}^{d-1})$ for $r>0$ and $p\geq 1$. Technically, they obtained the approximation error in $L^{p}$ norm including the case $p=\infty$. Two quantities including $\beta=\max\{1,(d+3+r)/(2(d-1))\}$ and $\tau\in[0,1]$ are involved in the rate. The paper Shen et al. (2022) established the convergence rates of the excess risk for classification with a class of convex loss functions. The target function of interests is also from the Sobolev space $W_{p}^{r}([0,1]^{d})$, and note that the feature domain is $[0,1]^{d}$. Under this setting, the most interesting aspect of this work is that they track the dependencies of the ambient dimension $d$ in the pre-factor hidden in the big-$\mathcal{O}$ notation. They showed the prefactor is polynomially dependent on $d$. This should be contrasted to the introduced works in the subsection 2.2, which have exponential dependences in the prefactor.
@@
The last three works in the above table give the dimension-free rates in the exponent of $n$ under different problem settings. The paper Zhou and Huo (2023) recently appeared on arXiv and studied the problem of learning binary classifiers through ReLU FNNs. An interesting aspect of this result is that unlike the other results Kim et al. (2021); Feng et al. (2021); Shen et al. (2022), the rate is both dimension and smoothness index free, despite the feature domain being unbounded. The main idea for getting this result is to leverage the fact that Gaussian distributions of features are analytic functions and have a fast decay rate, and this can be captured through ReLU FNNs. Another paper Ko et al. (2023) works on a similar problem but under a different setting. Notably, they are interested in the Barron Approximation (BA) space proposed in Caragea et al. (2023). Unlike the classical Barron space in Barron (1993), which is essentially a subset of a set of Lipschitz continuous functions, this space includes even discontinuous functions and hence is more general. The CAR assumption is imposed on the functions in the BA class. They proved that the rate $\mathcal{O}(n^{-\frac{q+1}{3(q+2)}})$ is achievable through ReLU FNNs for the estimation, and this rate is indeed minimax optimal. Note that the rate is slower than the parameteric rate $n^{-\frac{1}{2}}$ by noting the rate ranges from $n^{-\frac{1}{6}}$ to $n^{-\frac{1}{3}}$ as $q$ varies from 00 to $\infty$. This is attributed to the setting that the size of the distribution class is large.

### v3

@@ Appendix A Technical Results of Schmidt-Hieber (2020)
Several remarks follow in the sequel. First, the rate $\phi_{n}$ is dependent on the effective dimension $t_{i}$ which can be much less than the ambient input dimension $d_{0}$, and this implies that the neural net can circumvent the curse of dimensionality. Second, the depth $L$ should be chosen to scale with the sample size in a rate $\mathcal{O}(\log n)$. Third, the width $\mathbf{p}_{\text{max}}$ can be chosen to be independent of the smoothness indices. Lastly, the result is saying what is important for the statistical performance is not the sample size, but the amount of regularization. This is explicitly reflected through the number of active parameters $\mathcal{N}\asymp n\phi_{n}\log n\ll n$.
Appendix B Classification tasks via fully-connected networks
@@
We consider a binary classification problem via fully-connected networks. Classifiers built with neural networks handle large-scale high-dimensional data, such as facial images from computer vision extremely well, while traditional statistical methods often fail miserably. We will review papers that have attempted to provide theoretical explanations for such empirical successes in high-dimensional classification problems, beyond the existing statistical literature. This is a less studied area than nonparametric regression, although deep learning is often used for classification.
@@
Results on Neural Networks. The Table 1 summarizes the settings and results of the works that study the convergence rates of excess risks for binary classification problems of neural networks. Hereafter, $\alpha>0$ indicates the smoothness index of either target function class $(\Sigma)$ in CAR or boundary class $(\mathcal{G})$ in CAD, and $q>0$ denotes the term in the exponent in Margin Assumption (MA).
Reference Function Space Loss Condition Rate Kim et al. (2021) ReLU FNNs Hinge MA, CAD $\mathcal{O}\bigg{(}n^{-\frac{\alpha(q+1)}{\alpha(q+2)+(d-1)(q+1)}}\bigg{)}$ MA, CAR $\mathcal{O}\bigg{(}n^{-\frac{\alpha(q+1)}{\alpha(q+2)+d}}\bigg{)}$ Feng et al. (2021) ReLU CNNs Hinge CAR $\mathcal{O}\bigg{(}n^{-\frac{r}{2\beta(d-1)+r(2-\tau)}}\bigg{)}$ p-norm $\mathcal{O}\bigg{(}n^{-\frac{pr}{2(\beta+1)(d-1)+2pr(2-\tau)}}\bigg{)}$ 2-norm MA, CAR $\mathcal{O}\bigg{(}n^{-\frac{2rq}{(2+q)((\beta+1)(d-1)+2\tau)}}\bigg{)}$ Shen et al. (2022) Hinge MA, CAR $\mathcal{O}\bigg{(}n^{-\frac{r(q+1)}{d+2r(q+1)}}\bigg{)}$ Logistic CAR $\mathcal{O}\bigg{(}n^{-\frac{r}{2d+4r}}\bigg{)}$ Least Square $\mathcal{O}\bigg{(}n^{-\frac{4r}{3d+16r}}\bigg{)}$ Zhou and Huo (2023) ReLU FNNs Hinge MA: GMM $\mathcal{O}\bigg{(}n^{-\frac{q+1}{q+2}}\bigg{)}$ Ko et al. (2023) ReLU FNNs Logistic MA, CAR: BA $\mathcal{O}\bigg{(}n^{-\frac{q+1}{3(q+2)}}\bigg{)}$ Hu et al. (2020a) ReLU FNNs 0-1 Loss MA, CAD: Teacher-Student $\mathcal{O}\bigg{(}n^{-\frac{2}{3}}\bigg{)}$
Table 3: Comparison table on statistical convergence rates of neural networks in classification tasks.
@@
Two recent works studied the binary classification by ReLU convolutional neural networks Feng et al. (2021); Shen et al. (2022). The paper Feng et al. (2021) considered the $p$-norm loss $\phi(t):=\max\{0,1-t\}^{p}$ for $p\geq 1$, and the input data is supported on the sphere $\mathcal{S}^{d-1}$. In the paper, the approximation error bounds and excess risk bounds are derived under a varying power condition, and the target function class is the Sobolev space $W_{p}^{r}(\mathcal{S}^{d-1})$ for $r>0$ and $p\geq 1$. Technically, they obtained the approximation error in $L^{p}$ norm including the case $p=\infty$. Two quantities including $\beta=\max\{1,(d+3+r)/(2(d-1))\}$ and $\tau\in[0,1]$ are involved in the rate. The paper Shen et al. (2022) established the convergence rates of the excess risk for classification with a class of convex loss functions. The target function of interests is also from the Sobolev space $W_{p}^{r}([0,1]^{d})$, and note that the feature domain is $[0,1]^{d}$. Under this setting, the most interesting aspect of this work is that they track the dependencies of the ambient dimension $d$ in the pre-factor hidden in the big-$\mathcal{O}$ notation. They showed the prefactor is polynomially dependent on $d$.
@@
The last three works in the above table give the dimension-free rates in the exponent of $n$ under different problem settings. The paper Zhou and Huo (2023) studied the problem of learning binary classifiers through ReLU FNNs. An interesting aspect of this result is that unlike the other results Kim et al. (2021); Feng et al. (2021); Shen et al. (2022), the rate is both dimension and smoothness index free, despite the feature domain being unbounded. The main idea for getting this result is to leverage the fact that Gaussian distributions of features are analytic functions and have a fast decay rate, and this can be captured through ReLU FNNs. Another paper Ko et al. (2023) works on a similar problem but under a different setting. Notably, they are interested in the Barron Approximation (BA) space proposed in Caragea et al. (2023). Unlike the classical Barron space in Barron (1993), which is essentially a subset of a set of Lipschitz continuous functions, this space includes even discontinuous functions and hence is more general. The CAR assumption is imposed on the functions in the BA class. They proved that the rate $\mathcal{O}(n^{-\frac{q+1}{3(q+2)}})$ is achievable through ReLU FNNs for the estimation, and this rate is indeed minimax optimal. Note that the rate is slower than the parameteric rate $n^{-\frac{1}{2}}$ by noting the rate ranges from $n^{-\frac{1}{6}}$ to $n^{-\frac{1}{3}}$ as $q$ varies from 00 to $\infty$. This is attributed to the setting that the size of the distribution class is large.